if 6 and x have the same mean average as 2,4, and 24, what is the value of x
Answer by stanbon(48535)

if 6 and x have the same mean average as 2,4, and 24, what is the value of x
------------------------
The mean average of 2,4, and 24 is (2+4+24)/3 = 30/3 = 10
---------------------------
The mean average of 6 and x is (6+x)/2
-----------------------
EQUATION:
(6+x)/2 = 10
6+x = 20
x = 14
=============
Cheers,
Stan H.